### v3

# Removal and Selection: Improving RGB-Infrared Object Detection via Coarse-to-Fine Fusion

Abstract
Object detection in visible (RGB) and infrared (IR) images has been widely applied in recent years. Leveraging the complementary characteristics of RGB and IR images, the object detector provides reliable and robust object localization from day to night. Existing fusion strategies directly inject RGB and IR images into convolution neural networks, leading to inferior detection performance. Since the RGB and IR features have modality-specific noise, these strategies will worsen the fused features along with the propagation. Inspired by the mechanism of human brain processing multimodal information, this work introduces a new coarse-to-fine perspective to purify and fuse two modality features. Specifically, following this perspective, we design a Redundant Spectrum Removal module to coarsely remove interfering information within each modality and a Dynamic Feature Selection module to finely select the desired features for feature fusion. To verify the effectiveness of the coarse-to-fine fusion strategy, we construct a new object detector called Removal and Selection Detector (RSDet). Extensive experiments on three RGB-IR object detection datasets verify the superior performance of our method.
[…]
1 Introduction
With the development of various sensors, multimodal images are collected and applied in more and more application fields LI et al. (2021); Cao et al. (2015). Among them, visible (RGB) and infrared (IR) sensors are widely utilized due to their complementary imaging characteristics. Specifically, visible images can reflect rich texture and color information of the object but are greatly affected by low-light environments. In contrast, infrared images can reflect the thermal information of the object and can clearly reflect the outline of the object under poor illumination conditions, but lack the texture details of the object. In recent years, researchers have focused on the visible and infrared feature-level fusion  Yuan et al. (2022); Sun et al. (2022); Yuan and Wei (2023), which can help achieve better performance on downstream tasks (e.g., object detection).
Figure 1: Comparison between existing RGB-IR feature fusion structure and our proposed framework.
Figure 2: Effectiveness of our Coarse-to-Fine fusion. (a) is the current Halfway Fusion method, the directly extracted features are interfered with by the background information from the RGB image and suppress the final fused features, which will result in inferior detection results. (b) Our coarse-to-fine fusion can reduce the irrelevant information and select desired features for fusion, which achieves superior performance.
In RGB-IR object detection, the effective feature fusion method of RGB and IR images is crucial. Most existing RGB-IR object detection methods extract the modality-specific features from RGB and IR images independently, and then directly perform addition or concatenation operations on these features Liu et al. (2016a); Li et al. (2018); Cao et al. (2019), as shown in Figure 1(a). Without explicit cross-modal fusion, the “Late fusion” strategy is therefore limited in learning the complementary information, resulting in inferior performance. To further explore the optimal fusion strategies, many researchers have considered “Halfway fusion” strategy to design an interaction module between different modality features. For instance, Zhou et al. Zhou et al. (2020) construct the DMAF module to tap the difference between RGB and IR modalities which brings more useful information at the feature level. Xie et al. Xie et al. (2022) introduce a novel dynamic cross-modal module that aggregates local and global information from RGB and infrared modalities. Although these methods have achieved encouraging improvements, they ignore the negative impact of the modality-noise along with the propagation, which would fail to achieve complementary fusion.
Actually, when confronted with multi-modal information, our brains first establish rules to eliminate interfering information, and then meticulously select the desired information, which has been modeled in a cognition theory (“Attenuation Theory” Treisman (1964)). Such a way can be summarized as a coarse-to-fine process, which motivates us to introduce a new perspective to fuse RGB and IR features. As shown in Figure 1(c), we design a new fusion strategy called “Coarse-to-Fine Fusion” to achieve complementary feature fusion. “Coarse” indicates that our method begins with the filter out the interfering information and thus can coarsely remove the irrelevant spectrum. To this end, since the redundant information in an image also exists in its frequency spectrum, we propose a Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain. Specifically, we convert each source image into frequency space and introduce a dynamic filter to adaptively reduce irrelevant spectrum within RGB and IR modalities. As for “Fine”, it indicates that our fusion process conducts finely select features after the coarse removal. Therefore, we design a Dynamic Feature Selection (DFS) module to meticulously select the desired features between RGB and IR modalities. In this module, we gate different scale features required for object detection by exploring a mixture of scale-aware experts. Figure 2 gives an example of Coarse-to-Fine fusion strategy. To evaluate the effectiveness of the coarse-to-fine strategy, we build a novel framework for RGB-IR object detection called Removal and Selection Detector (RSDet).
In summary, this paper has the following contributions:
  • We present a new coarse-to-fine perspective to fuse RGB and IR features. Inspired by the mechanism of human brain processing multimodal information, we coarsely remove the interfering information and finely select desired features for fusion.
  • Following the coarse-to-fine fusion strategy, we propose a Redundant Spectrum Removal module, which introduces a dynamic spectrum filter to adaptively reduce irrelevant information in the frequency domain. Besides, we also design a Dynamic Feature Selection module, which utilize a mixture of scale-aware experts to gate different scale features for final fusion.
  • To verify the effectiveness of coarse-to-fine strategy, we build a novel object detector based on it for RGB-IR object detection. Extensive experiments on three public RGB-IR object detection datasets demonstrate that our proposed method achieves state-of-the-art performance.
2 Related Works
2.1 Multispectral Object Detection
In recent years, thanks to the development of deep learning technology and several visible and infrared datasets being proposed Hwang et al. (2015); Zhang et al. (2020); Jia et al. (2021), the multispectral object detection task has gradually attracted attention. To fully explore the effective information between visible and infrared images, some researchers focus on the complementarity between the two modalities starting from the illumination conditions. Guan et al. Guan et al. (2019) and Li et al. Li et al. (2019) first propose the illumination-aware modules to allow the object detectors to adjust the fusion weight based on the predicted illumination conditions. Moreover, Zhou  et al. Zhou et al. (2020) analyze and address the modality imbalance problems by designing two feature fusion modules called DMAF and IAFA. Recently, an MSR memory module Kim et al. (2022) was introduced to enhance the visual representation of the single modality features by recalling the multispectral modality features, which enables the detector to encode more discriminative features. Yuan et al. Yuan et al. (2024) propose a transformer-based RGB-IR object detector to further improve the object detector performance. Different from these works, we present a new coarse-to-fine fusion strategy to fuse RGB and IR features for multispectral object detection task.
Figure 3: The overall architecture of RSDet. The entire network follows the “Coarse-to-Fine” fusion and comprises a two-stream backbone network taking RGB and IR inputs, a Redundant Spectrum Removal (RSR) module, and a Dynamic Feature Selection (DFS) module.
2.2 Shared-Specific Representation Learning
Shared and specific representation learning is first explored in the Domain Separation Network Bousmalis et al. (2016) for unsupervised domain adaptation. It uses a shared-weight encoder to capture shared features and a private encoder to capture domain–specific features. Sanchez et al. Sanchez et al. (2020) explored further shared and specific feature disentanglement representation, and found it is useful to perform downstream tasks such as image classification and retrieval. Recently, van et al. van Amsterdam et al. (2023) improved the performance of action segmentation by disentangling the latent features into shared and modality-specific components. Furthermore, Wang et al. Wang et al. (2023) proposed the ShaSpec model handled missing modalities problems. Shared-specific representation learning has shown great performance and effectiveness in feature learning. However, few RGB-IR object detection models explicitly exploit shared-specific representation. In this paper, we introduce shared-specific representation learning between RGB and IR modality features to implement our coarse-to-fine fusion strategy.
2.3 Mixture of Experts
The Mixture-of-Experts (MoE) model Jacobs et al. (1991); Jordan and Jacobs (1994) has demonstrated the ability to dynamically adapt its structure based on varying input conditions. Several studies have been dedicated to the theoretical exploration of MoE Lepikhin et al. (2021); Nie et al. (2022); Zhou et al. (2022), focusing on enhancing model quality, efficient training and router mechanisms. Besides, some researchers also concentrated on leveraging the MoE model for specific tasks. For example, Gross et al. Gross et al. (2017) observed that a hard mixture-of-experts model can be efficiently trained to good effect on large scale multi-label prediction tasks. Cao et al. Cao et al. (2023a) proposed a mixture of local-to-global experts (MoE-Fusion) mechanism by integrating MoE structure into image fusion tasks. Chen et al. Chen et al. (2023) addressed the multi-task learning challenge by implementing a cooperative and specialized mechanism among experts. In our proposed method, the MoE model is integrated into the multispectral object detection task. Specifically, we establish multi-scale experts to gate modality features at different scales and leverage its dynamic fusion mechanism to facilitate complementary fusion.
3 Method
3.1 Coarse-to-Fine Fusion
The design of the “Coarse-to-Fine Fusion” strategy is shown in Figure 3, “Coarse” corresponding to the Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain and ‘Fine’ corresponding to the Dynamic Feature Selection (DFS) module to meticulously select the desired features between RGB-IR modalities. Since the two modality features are usually intersection, we introduce disentangled representation learning Sanchez et al. (2020) to purify and decouple them for complementary fusion. Thus, we integrate above two modules into the shared-specific structures to implement our coarse-to-fine fusion. Specifically, we first perform the RSR module to remove interfering information. Then, we introduce the shared-specific structure to extract the two modality-specific multiscale features $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$, which uses ResNet50 He et al. (2016) as the backbone network. As for the shared features $C_{\text{sha}}$, we also employ several Resblocks as the feature extractor. After that, these different scale features $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$ are dynamically aggregated by exploring a mixture of scale-aware experts and obtain the final specific feature $C_{\text{spe}}$ by DFS module. Finally, the specific feature $C_{\text{spe}}$ and the shared feature $C_{\text{sha}}$ are added together for the final object detection task, which can be expressed as :
$C=C_{\text{sha}}+C_{\text{spe}}.$ (1)
In the next sections, we will further introduce the proposed RSR module and DFS module.
3.2 Redundant Spectrum Removal
For the “Coarse Removal”, since the frequency domain obeys the nature of global modeling, we propose a Redundant Spectrum Removal (RSR) module to filter out coarsely in the frequency domain. Specifically, we convert each source image into frequency space and predict a dynamic filter to adaptively reduce irrelevant spectrum within RGB and IR modalities. Specifically, the paired RGB image $V\in\mathbb{R}^{H\times W\times 3}$ and IR image $I\in\mathbb{R}^{H\times W}$ are token as the RSR input. They are first subjected to the $\operatorname{DFT}(\cdot)$ operation and get the frequency domain image $F_{I}(u,v)$ and $F_{V}(u,v)$:
$F_{I}(u,v)=\operatorname{DFT}(I),~{}F_{V}(u,v)=\operatorname{DFT}(V).$ (2)
The filter prediction network is conducted to dynamically generate the redundant spectrum filter, which is illustrated in Figure 3. For each modality image, we perform an embedding operation on the entire image get a feature vector, and reshape it to match the size of the original image:
${M}_{l_{I}}=\operatorname{Embed}(I),$ ${M}_{l_{V}}=\operatorname{Embed}(V),$ (3)
$M_{m_{I}}=\operatorname{Reshape}(M_{l_{I}}),$ $M_{m_{V}}=\operatorname{Reshape}(M_{l_{V}}).$
Then, we perform the $\operatorname{top}K$ operation on $M_{m}$ to get the binary filters $H_{I}(u,v)$ and $H_{V}(u,v)$. In specific, the positions corresponding to the $\operatorname{top}K$ values are set to 1 and the others to 0:
$H_{I}(u,v)=\operatorname{top}K(M_{m_{I}}),H_{V}(u,v)=\operatorname{top}K(M_{m_{V}}).$ (4)
Finally, we apply $H(u,v)$ to filter the frequency domain image $F(u,v)$, like $F_{I}^{{}^{\prime}}(u,v)=F_{I}(u,v)\otimes H_{I}(u,v)$ and $F_{V}^{{}^{\prime}}(u,v)=F_{V}(u,v)\otimes H_{V}(u,v)$ and subject the Inverse Discrete Fourier Transform $\operatorname{IDFT}(\cdot)$ to transform image back to the spatial domain. Thus, we obtain the final irrelevant spectrum removed images $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$.
$I^{{}^{\prime}}=\operatorname{IDFT}(F_{I}^{{}^{\prime}}(u,v)),~{}V^{{}^{\prime}}=\operatorname{IDFT}(F_{V}^{{}^{\prime}}(u,v)).$ (5)
3.3 Dynamic Feature Selection
For the “Fine Selection”, we implement the dynamic modality feature selection in the fusion process by employing a mixture of scale-aware experts to gate multi-scale features. As shown in Figure 3, we design a dedicated expert for each scale modality-specific feature. Then, we aggregate these features by using the gating network to predict a set of dynamic weights. Specifically, after obtaining different scale features $C_{\text{I-spe}}^{i}$ and $C_{\text{V-spe}}^{i}$ through the feature extraction for $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$, we first make $C_{\text{I-spe}}^{i}$ and $C_{\text{V-spe}}^{i}$ go through the average pooling operations, and flattened them into one-dimension vector $X_{I}^{i}\in\mathbb{R}^{M}$ and $X_{V}^{i}\in\mathbb{R}^{M}$ to predict the weights $w_{I}^{i}$ and $w_{V}^{i}$ of the gating network $G$. It can be formulated as follows:
$w_{I}^{i},w_{V}^{i}=G\left(X_{I}^{i},X_{V}^{i}\right)=\operatorname{Softmax}\left(\left[X_{I}^{i},X_{V}^{i}\right]\cdot W_{\text{}}\right),$ (6)
where the $W\in\mathbb{R}^{M\times N}$ is the learnable weight matrix normalized through the softmax operation and $i$ is the index of experts. After that, the weights $w_{I}$ and $w_{V}$ are converted to gating via the Router $R$, preserving the desired features between the two modalities for fusion at different scales. Consequently, the Router $R$ can be formulated as follows:
$\begin{gathered}(r_{I}^{i},r_{V}^{i})=R(w_{I}^{i},w_{V}^{i})=\begin{cases}(1,1),&w_{I}^{i},w_{V}^{i}\geq t\\ (1,0),&w_{I}^{i}\geq t,w_{V}^{i}\leq t,\\ (0,1),&w_{I}^{i}\leq t,w_{V}^{i}\geq t\end{cases}\end{gathered}$ (7)
where $t$ is a threshold. Then $N$ scale-aware expert networks $\mathcal{E}_{I}^{i}$ and $\mathcal{E}_{V}^{i}$ are utilized to further extract modality-specific features. The formalization is as follows:
$C_{I}^{i}=\mathcal{E}_{I}^{i}(x_{I}^{i}\cdot r_{I}^{i}),~{}C_{V}^{i}=\mathcal{E}_{V}^{i}(x_{V}^{i}\cdot r_{V}^{i}),$ (8)
where $x_{I}^{i}$ and $x_{V}^{i}$ are the multiscale features of different modality input $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$. The detailed structure of each scale-aware expert $\mathcal{E}$ is the same and mainly consists of two convolution blocks. After obtaining the output results from expert models at different scales, we perform dynamic weighted summation and concatenate them together to get the ultimate multi-modal specific feature $C_{\text{spe}}$:
$C_{\text{spe}}=\bigcup_{i=1}^{n}\left(w_{I}^{i}C_{I}^{i}+w_{V}^{i}C_{V}^{i}\right).$ (9)
Table 1: Detection results ($\text{MR}^{\text{-2}}$, in%) under ‘All-dataset’ settings of different pedestrian distances, occlusion levels, and light conditions (Day and Night) on the KAIST dataset. The pedestrian distances consist of ‘Near’ (115 $\leq$ height), ‘Medium’ (45 $\leq$ height $<$ 115) and ‘Far’ (1 $\leq$ height $<$ 45), while occlusion levels consist of ‘None’ (never occluded), ‘Partial’ (occluded to some extent up to half) and ‘Heavy’ (mostly occluded). IoU $=0.5$ is set for evaluation. The best results are highlighted in red and the second-place are highlighted in blue.
Methods Backbone Near Medium Far None Partial Heavy Day Night All
ACF Hwang et al. (2015) VGG16 28.74 53.67 88.20 62.94 81.40 88.08 64.31 75.06 67.74
Halfway Fusion Liu et al. (2016a) VGG16 8.13 30.34 75.70 43.13 65.21 74.36 47.58 52.35 49.18
FusionRPN+BF Konig et al. (2017) VGG16 0.04 30.87 88.86 47.45 56.10 72.20 52.33 51.09 51.70
IAF R-CNN Li et al. (2019) VGG16 0.96 25.54 77.84 40.17 48.40 69.76 42.46 47.70 44.23
IATDNN+IASS Guan et al. (2019) VGG16 0.04 28.55 83.42 45.43 46.25 64.57 49.02 49.37 48.96
CIAN Zhang et al. (2019a) VGG16 3.71 19.04 55.82 30.31 41.57 62.48 36.02 32.38 35.53
MSDS-R-CNN Li et al. (2018) VGG16 1.29 16.19 63.73 29.86 38.71 63.37 32.06 38.83 34.15
AR-CNN Zhang et al. (2019b) VGG16 0.00 16.08 69.00 31.40 38.63 55.73 34.36 36.12 34.95
MBNet Zhou et al. (2020) ResNet50 0.00 16.07 55.99 27.74 35.43 59.14 32.37 30.95 31.87
TSFADet Yuan et al. (2022) ResNet50 0.00 15.99 50.71 25.63 37.29 65.67 31.76 27.44 30.74
CMPD Li et al. (2022) ResNet50 0.00 12.99 51.22 24.04 33.88 59.37 28.30 30.56 28.98
CAGTDet Yuan et al. (2024) ResNet50 0.00 14.00 49.40 24.48 33.20 59.35 28.79 27.73 28.96
RSDet (Ours) ResNet50 0.00 12.13 39.80 20.49 33.25 57.60 25.83 26.48 26.02
3.4 Removal and Selection Detector (RSDet)
To evaluate the effectiveness of our Coarse-to-Fine fusion strategy, we embed it into an existing object detection framework. In specific, we utilize a two-stage detector Faster R-CNN Ren et al. (2017) as our base model and replace its backbone with our strategy to construct a new object detector called RSDet. Other modules such as Region Proposal Network (RPN) and R-CNN head remain unchanged.
Loss functions. To extract tyhe shared and specific feature from the images $I^{{}^{\prime}}$ and $V^{{}^{\prime}}$, we maximize the mutual information between $C_{\text{I-spe}}$ and $C_{\text{V-spe}}$ with $C_{\text{sha}}$ The mutual information can serve as the deep supervising loss function $\mathcal{L}_{\text{I-spe}}$ and $\mathcal{L}_{\text{V-spe}}$ to guide the shared-specific features learning. The definitions are as follows:
$\mathcal{L}_{\text{I-spe}}=\operatorname{MI}(C_{\text{sha}},C_{\text{I-spe}}),$ (10)
$\mathcal{L}_{\text{V-spe}}=\operatorname{MI}(C_{\text{sha}},C_{\text{V-spe}}),$ (11)
where $\operatorname{MI}$ represents the operation of mutual information. We use cross-entropy ($\operatorname{CE}$) and KL-divergence ($\operatorname{KL}$) to calculate the mutual information Sanghi (2020); Tschannen et al. (2019) between different features in the latent space.
$\max\operatorname{MI}(x,y)\Rightarrow\max\{\operatorname{CE}(x,y)-\operatorname{KL}(x||y)$ (12)
$+\operatorname{CE}(y,x)-\operatorname{KL}(y||x)\}.$
For detection loss, we also use the detection loss $\mathcal{L}_{\text{rpn}}$, $\mathcal{L}_{\text{reg}}$ and $\mathcal{L}_{\text{cls}}$ same as the Faster R-CNN Ren et al. (2017) to supervise the whole RSDet network training. Thus, the total loss is the sum of these individual losses:
$\mathcal{L}=\gamma(\mathcal{L}_{\text{I-spe}}+\mathcal{L}_{\text{V-spe}})+\mathcal{L}_{\text{rpn}}+\mathcal{L}_{\text{reg}}+\mathcal{L}_{\text{cls}},$ (13)
where $\gamma=0.001$ is the coefficient used to strike a balance between the different loss functions.
4 Experiments
4.1 Datasets
KAIST Dataset. KAIST Hwang et al. (2015) is a public aligned multispectral pedestrian detection dataset. Due to the problematic annotations in the original dataset, our framework is trained on 8963 image pairs with new annotations Zhang et al. (2019b) and evaluated on 2252 image pairs with the improved annotations Liu et al. (2016a). We perform comparison experiments under ‘All-dataset’ settings Hwang et al. (2015) and evaluate six subsets in terms of pedestrian distance and occlusion level.
FLIR Dataset. FLIR is a paired visible and infrared object detection dataset including daytime and night scenes. Since the images are misaligned in the original dataset, we use the FLIR-aligned dataset Zhang et al. (2020) in our experiment. It has 5142 aligned RGB-IR image pairs, of which 4129 are used for training and 1013 for testing.
LLVIP Dataset. LLVIP Jia et al. (2021) is a strictly aligned visible and infrared object detection dataset for low-light vision. It is collected in low-light environments most of which were captured in very dark scenes. It contains 15488 aligned RGB-IR image pairs, of which 12025 images are used for training and 3463 images for testing.
4.2 Evaluation Metrics
log-average Miss Rate ($\text{MR}^{\text{-2}}$). For KAIST dataset, we employ the standard KAIST evaluation Hwang et al. (2015): log-average Miss Rate (MR) over False Positive Per Image (FPPI) range of $[10^{-2},10^{0}]$ (also denoted as $\text{MR}^{\text{-2}}$). The lower score indicates better performance.
mean Average Precision (mAP). For FLIR and LLVIP datasets, we employ the commonly used object detection metric $\text{mAP}_{50}$ and mAP. The $\text{mAP}_{50}$ metric represents the mean AP under IoU 0.50 and The mAP metric represents the mean AP under IoU ranges from 0.50 to 0.95 with a stride of 0.05. The higher value of two metrics means better performance.
Table 2: Camparison the performance (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in red and the second-place are highlighted in blue.
FLIR LLVIP
[…]
SSD Liu et al. (2016b) VGG16 65.5 29.6 90.2 53.5
RetinaNet Lin et al. (2017) ResNet50 66.1 31.5 94.8 55.1
Cascade R-CNN Cai and Vasconcelos (2019) ResNet50 71.0 34.7 95.0 56.8
Faster R-CNN Ren et al. (2017) ResNet50 74.4 37.6 94.6 54.5
DDQ-DETR Zhang et al. (2023) ResNet50 73.9 37.1 93.9 58.6 IR
SSD Liu et al. (2016b) VGG16 52.2 21.8 82.6 39.8
RetinaNet Lin et al. (2017) ResNet50 51.2 21.9 88.0 42.8
Cascade R-CNN Cai and Vasconcelos (2019) ResNet50 56.0 24.7 88.3 47.0
Faster R-CNN Ren et al. (2017) ResNet50 64.9 28.9 87.0 45.1
DDQ-DETR Zhang et al. (2023) ResNet50 64.9 30.9 86.1 46.7 RGB
Halfway fusion Liu et al. (2016a) VGG16 71.5 35.8 91.4 55.1
GAFF Zhang et al. (2021) ResNet18 74.6 37.4 94.0 55.8
ProbEn Chen et al. (2022) ResNet50 75.5 37.9 93.4 51.5
LGADet Zuo et al. (2023) ResNet50 74.5 - - -
CSAA Cao et al. (2023b) ResNet50 79.2 41.3 94.3 59.2
RSDet (Ours) ResNet50 81.1 41.4 95.8 61.3 RGB+IR
4.3 Implementation Details
All the experiments are implemented in the mmdetection Chen et al. (2019) toolbox and conducted on a single GPU NVIDIA GeForce RTX 3090. We use ResNet-50 as the backbone. During training, the batch size is set to 4 and the SGD optimizer is employed with a momentum of 0.9 and the weight decay of $1\times 10^{-4}$. The input image resolution is $640\times 512$ for FLIR and KAIST and $1280\times 1024$ for LLVIP. All experiments are trained for 12 epochs with an initial learning rate of $1\times 10^{-2}$. We use random flipping with a probability of 0.5 to increase input diversity.
Figure 4: The FPPI-MR curves comparisons with the state-of-the-art methods on the KAIST dataset under the ‘All-dataset’ settings Hwang et al. (2015).
4.4 Comparison with State-of-the-Art Methods
KAIST Dataset. The experiments are conducted on the KAIST under ‘All-dataset’ settings Hwang et al. (2015); Li et al. (2018). We compare the proposed RSDet with twelve SOTA multispectral object detection methods on the KAIST dataset. As shown in Table 1, the RSDet performs best on the ‘All’, ‘Day’, and ‘Night’ conditions and the other four of six subsets (‘Near’, ‘Medium’, ‘Far’, ‘None’). And rank second in two subsets (Partial and Heavy). Specifically, the RSDet takes the lead in the subsets of different scales (‘Near’, ‘Medium’, ‘Far’). Especially, in the ‘Far’ subsets, the RSDet surpasses the second-best by about 10%, which is attributed to the mixture of scale-aware experts in the DFS module. Besides, on the whole test dataset (‘All’), the RSDet surpasses the second-best by about 3%, which indicates the satisfactory accuracy of RSDet. The log-average Miss Rate ($\text{MR}^{\text{-2}}$) over the False Positive Per Image (FPPI) curve of all detectors is also shown in Figure 4. The above results prove the superiority of our method.
FLIR and LLVIP Dataset. The quantitative results of the different methods on the FLIR and LLVIP datasets are shown in Table 2. We compare the proposed RSDet with five SOTA multispectral object detection methods and five single-modality object detection methods. From Table 2, it can be observed that the experimental results of fusion methods are generally better than those of single-modal methods in the FLIR dataset. However, due to the limited light conditions, the RGB features interfere with the IR features in the LLVIP dataset, resulting in the detection results of the multispectral detector being inferior to the IR modality. However, our method effectively addresses this issue with the well-designed coarse-to-fine fusion strategy and achieves state-of-the-art performance on the FLIR and LLVIP datasets.
Table 3: Ablation study on each module result (mAP, in%) on the FLIR and LLVIP dataset. The best results are highlighted in bold.
FLIR LLVIP
RSR DFS $\text{mAP}_{50}$ mAP $\text{mAP}_{50}$ mAP
75.2 37.2 95.0 55.5
76.3 38.4 95.0 57.1
81.1 41.4 95.8 61.3
Figure 5: Visualization of DFS module feature fusion results on the FLIR dataset. To facilitate a clearer observation, we overlaid the features onto the original RGB or IR image.
Figure 6: Ablation study on different value $K$ in RSR module on the FLIR and LLVIP dataset.
4.5 Ablation Study
Ablation on each component. To verify the effectiveness of the RSR and DFS modules, we ablate each module from RSDet. As shown in Table 3, after adding the RSR module the performance increases 1.2% mAP. on the FLIR dataset and 1.6% mAP on the LLVIP dataset respectively. Further adding the DFS module, the metrics show a significant increase on the FLIR dataset, the $\text{mAP}_{50}$ increase 4.8%, and the mAP increase 3.0%. On the LLVIP dataset, the experiment result also exhibits similar trends, especially in the mAP metric. The above results indicate the effectiveness of these modules.
Ablation on different K.  We further study the impact of different values of $K$ in the RSR module. We conduct the experiments by modifying the value of $K$ ranging from 100 to 400 on the FLIR and LLVIP datasets. As shown in Figure 6, our detector achieves the highest detection performance when $K=320$. If the value of $K$ is increased or decreased, the performance of the detector will decrease, which also proves the negative impact of the redundant spectrum for the detector.
4.6 Visualization and Discussion
RSR module. We conduct tSNE visualizations Van der Maaten and Hinton (2008) of shared $C_{\text{sha}}$ and specific features ($C_{\text{I-spe}}$ and $C_{\text{V-spe}}$) on FLIR dataset. As shown in Figure 7, we have observed that without the RSR module, there is still a considerable mix of feature points in the shared and specific features, leading to difficulty in selecting the desired specific features in DFS. However, after adding the RSR module, we have noted a significant decrease in the number of mixed features, which verifies that removing redundant spectra is beneficial to feature disentanglement and, thus more effective in obtaining fused features in the DFS module.
Figure 7: tSNE visualization Van der Maaten and Hinton (2008) of the modality-specific and shared features. (a) ‘w/o RSR’ represents without RSR module. (b) ‘w RSR’ represents without RSR module.
[…]
5 Conclusion
In this paper, we presented a new coarse-to-fine perspective to fuse visible and infrared modality features. Specifically, a Redundant Spectrum Removal (RSR) module is first designed to coarsely filter out the irrelevant spectrum, and then a Dynamic Feature Selection (DFS) module is proposed to finely select the desired features for the final fusion process. we constructed a new object detector called Removal and Selection Detector (RSDet) to evaluate its effectiveness and versatility. Extensive experiments on three public RGB-IR detection datasets demonstrated that our method can effectively facilitate complementary fusion and achieve state-of-the-art performance. We believe that our method can be applied to various studies in the RGB-IR feature fusion tasks.